### v1

@@ B.1 Proof of Proposition˜1
For any matrices $A,B\in\real{n\times n}$, the singular-value perturbation bounds [41, Corollary 7.3.5] state that
$\begin{split}&|\sigma_{\max}(A)-\sigma_{\max}(B)|\leq\|A-B\|_{2},\\ &|\sigma_{\min}(B)-\sigma_{\min}(A)|\leq\|A-B\|_{2}.\end{split}$ (38)
Apply (LABEL:eqn:sig_per) with $A=X^{1/2}$ and $B=\hat{X}^{1/2}U$, and combine with (37) to obtain
$\begin{split}&|\sigma_{\max}(X^{1/2})-\sigma_{\max}(\hat{X}^{1/2}U)|\leq\theta,\\ &|\sigma_{\min}(\hat{X}^{1/2}U)-\sigma_{\min}(X^{1/2})|\leq\theta\end{split}$
Because $U$ is orthogonal, $X\succeq 0$ and $\hat{X}\succeq 0$, we have
[…]
Substituting these into the inequalities above yields
$\sqrt{\lambda_{\max}(X)}=\sigma_{\max}(X^{1/2})\leq\sqrt{\lambda_{\max}(\hat{X})}+\theta,$
$\sqrt{\lambda_{\min}(X)}=\sigma_{\min}(X^{1/2})\geq\sqrt{\lambda_{\min}(\hat{X})}-\theta$
Squaring both sides and clipping the lower bound at zero yields the claim. ∎
@@ B.2 Proof of Corollary˜2
The lower bounds follow directly from the constraints
$\Sigma_{v,t}\succeq\lambda_{\min}(\hat{\Sigma}_{v,t})I_{n_{y}},\quad\Sigma_{w,t-1}\succeq\lambda_{\min}(\hat{\Sigma}_{w,t-1})I_{n_{x}},\quad\Sigma_{x,0}^{-}\succeq\lambda_{\min}(\hat{\Sigma}_{x,0}^{-})I_{n_{x}},$
which are imposed in (4) and (5). Thus, any maximizers $\Sigma_{v,t}^{*}$, $\Sigma_{w,t-1}^{*}$, and $\Sigma_{x,0}^{-,*}$ necessarily satisfy these inequalities, giving the claimed lower bounds. ∎
@@ B.3 Proof of Theorem˜2
Proof.
Define the posterior and prediction maps
$\begin{split}&\Psi(\Sigma_{x,t}^{-},\Sigma_{v,t})\coloneqq\Sigma_{x,t}^{-}-\Sigma_{x}C_{t}^{\top}(C_{t}\Sigma_{x,t}^{-}C_{t}^{\top}+\Sigma_{v,t})^{-1}C_{t}\Sigma_{x,t}^{-}\\ &\Phi(\Sigma_{x,t},\Sigma_{w,t})\coloneqq A_{t}\Sigma_{x,t}A_{t}^{\top}+\Sigma_{w,t}.\end{split}$
The DRKF recursion is obtained by evaluating these maps at the least-favorable noise covariances:
$\Sigma_{x,t}=\Psi(\Sigma_{x,t}^{-},\Sigma_{v,t}^{*})\quad\text{and}\quad\Sigma_{x,t+1}^{-}=\Phi(\Sigma_{x,t},\Sigma_{w,t}^{*}).$
Using the identity $\Psi(\Sigma_{x,t}^{-},\Sigma_{v,t})=((\Sigma_{x,t}^{-})^{-1}+C_{t}^{\top}\Sigma_{v,t}^{-1}C_{t})^{-1}$ and the fact that $P_{1}\succeq P_{2}$ implies $P_{1}^{-1}\preceq P_{2}^{-1}$, the map $\Psi$ is monotone increasing in each argument. The map $\Phi$ is also monotone since congruence with $A_{t}$ and addition of a PSD matrix preserve the Loewner order.
Corollary˜2, together with the monotonicity of the posterior and prediction maps $\Psi$ and $\Phi$, implies by induction that the KF driven by $(\underline{\lambda}_{w,t}I_{n_{x}},\underline{\lambda}_{v,t}I_{n_{y}})$ produces covariances no larger than those of the DRKF, and the KF driven by $(\overline{\lambda}_{w,t}I_{n_{x}},\overline{\lambda}_{v,t}I_{n_{y}})$ produces covariances no smaller. ∎
@@ Appendix C Proofs for Convergence
C.1 Details of the Downsampled System
For the convergence analysis, we define a downsampled state as $x_{t}^{d}=x_{tN}$ for some $N\in\mathbb{N}$. Then, the downsampled dynamics take the form
$\begin{split}x_{t+1}^{d}&=A^{N}x_{t}^{d}+\mathcal{R}_{N}\bm{w}_{t}^{N}\\ y_{t}^{N}&=\mathcal{O}_{N}x_{t}^{d}+\mathcal{D}_{N}\bm{v}_{t}^{N}+\mathcal{H}_{N}\bm{w}_{t}^{N}\\ \bm{0}&=\mathcal{O}_{N}^{R}x_{t}^{d}+\bm{u}_{t}^{N}+\mathcal{L}_{N}\bm{w}_{t}^{N},\end{split}$
where the block matrices collecting process and measurement noise are
$\begin{split}\mathcal{R}_{N}&\coloneqq\left[\hat{\Sigma}_{w}^{\frac{1}{2}},\,A\hat{\Sigma}_{w}^{\frac{1}{2}},\,\ldots,\,A^{N-1}\hat{\Sigma}_{w}^{\frac{1}{2}}\right]\\ \mathcal{O}_{N}&\coloneqq\left[(CA^{N-1})^{\top},\,\ldots,\,(CA)^{\top},\,C^{\top}\right]^{\top}\\ \mathcal{O}_{N}^{R}&\coloneqq\left[(A^{N-1})^{\top},\,\ldots,\,(A)^{\top},\,I_{n_{x}}\right]^{\top}\\ \mathcal{D}_{N}&\coloneqq I_{N}\otimes\hat{\Sigma}_{v}^{\frac{1}{2}}.\end{split}$
Here, $\mathcal{R}_{N}$ and $\mathcal{O}_{N}$ represent the $N$-block controllability and observability matrices. The block Hankel matrices $\mathcal{L}_{N}$ and $\mathcal{H}_{N}\coloneqq(I_{N}\otimes C)\mathcal{L}_{N}$ encode cross-correlations between observations and noise terms, with entries of $\mathcal{L}_{N}$ defined as $\mathcal{L}_{N}]_{ij}=A^{j-i-1}\hat{\Sigma}_{w}^{\frac{1}{2}}$ for $j>i$ and zero otherwise.
Adopting the procedure in [36] and denoting the downsampled prior covariance matrix by $\Sigma_{x,t}^{-,d}:=\Sigma_{x,tN}^{-}$, the DR Riccati recursion associated with the downsampled model is given by (23), where
$r_{\mathbb{D},t}^{d}(\Sigma_{x,t}^{-,d})\coloneqq\alpha_{N,t}\left[(\Sigma_{x,t}^{-,d})^{-1}+\Omega_{\bar{\Phi}_{N,t}}\right]^{-1}\alpha_{N,t}+W_{\bar{\Phi}_{N,t}},$
with auxiliary matrices given by
$\alpha_{N,t}$ $\coloneqq A^{N}-\mathcal{R}_{N}(\mathcal{H}_{N}^{\top}\mathcal{K}_{{\bar{\Phi}}_{N,t}}^{-1}\mathcal{O}_{N}+\mathcal{L}_{N}^{\top}\mathcal{K}_{{\bar{\Phi}}_{N,t}}^{-1}\mathcal{O}_{N}^{R})$
$W_{\bar{\Phi}_{N,t}}$ $\coloneqq\mathcal{R}_{N}\mathcal{Q}_{\bar{\Phi}_{N,t}}\mathcal{R}_{N}^{\top}$
$\mathcal{Q}_{\bar{\Phi}_{N,t}}$ $\coloneqq\left[I_{Nn_{x}}+\mathcal{H}_{N}^{\top}(\mathcal{D}_{N}\mathcal{D}_{N}^{\top})^{-1}\mathcal{H}_{N}-\mathcal{L}_{N}^{\top}\bar{\Phi}_{N,t}\mathcal{L}_{N}\right]^{-1}$
$\Omega_{\bar{\Phi}_{N,t}}$ $\coloneqq\Omega_{N}+\mathcal{J}_{N}^{\top}S_{\bar{\Phi}_{N,t}}^{-1}\mathcal{J}_{N}$
$\mathcal{J}_{N}$ $\coloneqq\mathcal{O}_{N}^{R}-\mathcal{L}_{N}\mathcal{H}_{N}^{\top}\left[\mathcal{D}_{N}\mathcal{D}_{N}^{\top}+\mathcal{H}_{N}\mathcal{H}_{N}^{\top}\right]^{-1}\mathcal{O}_{N}$
$\Omega_{N}$ $\coloneqq\mathcal{O}_{N}^{\top}(\mathcal{D}_{N}\mathcal{D}_{N}^{\top}+\mathcal{H}_{N}\mathcal{H}_{N}^{\top})^{-1}\mathcal{O}_{N}$
$S_{\bar{\Phi}_{N,t}}$ $\coloneqq\mathcal{L}_{N}(I_{Nn_{x}}+\mathcal{H}_{N}^{\top}(\mathcal{D}_{N}\mathcal{D}_{N}^{\top})^{-1}\mathcal{H}_{N})^{-1}\mathcal{L}_{N}^{\top}-\bar{\Phi}_{N,t}^{-1}$
$\mathcal{K}_{\bar{\Phi}_{N,t}}$ $\coloneqq\begin{bmatrix}\mathcal{D}_{N}\mathcal{D}_{N}^{\top}+\mathcal{H}_{N}\mathcal{H}_{N}^{\top}&\mathcal{H}_{N}\mathcal{L}_{N}^{\top}\\ \mathcal{L}_{N}\mathcal{H}_{N}^{\top}&\mathcal{L}_{N}\mathcal{L}_{N}^{\top}-\bar{\Phi}_{N,t}^{-1}\end{bmatrix}$.
@@ C.2 Proof of Proposition˜2
Proof.
The proof follows the argument in [36, Proposition 3.1]. The downsampled DR Riccati mapping in (23) has the same structure as the robust Riccati mapping analyzed in [42, Theorem 5.3]. According to this result, if the matrices $\Sigma_{x,t}^{-,d},\Omega_{\bar{\Phi}_{N,t}}$, and $W_{\bar{\Phi}_{N,t}}$ are positive definite, then $r_{\mathbb{D},t}^{d}$ is a contraction mapping.
By Assumption˜1, $\hat{\Sigma}_{w}\succ 0$, which ensures that $\Sigma_{x,t}^{-,d}\succ 0$. Moreover, $\mathcal{Q}_{N,t}\succ 0$ whenever $0\preceq\bar{\Phi}_{N,t}\prec\tilde{\phi}_{N}I_{Nn_{x}}$, which, under the controllability assumption, is sufficient for the positive definiteness of $W_{\bar{\Phi}_{N,t}}$ for $N\geq n_{x}$.
Similarly, $S_{\bar{\Phi}_{N,t}}\prec 0$ and the map $\bar{\Phi}_{N,t}\mapsto\Omega_{\bar{\Phi}_{N,t}}$ is monotonically non-increasing for $0\prec\bar{\Phi}_{N,t}\prec\tilde{\phi}_{N}I_{Nn_{x}}$. Under the observability assumption, the matrix $\Omega_{0}=\Omega_{N}$ is positive definite for $N\geq n_{x}$. Hence, by continuity, there exists $\phi_{N}\in(0,\tilde{\phi}_{N})$ such that $\Omega_{\bar{\Phi}_{N,t}}\succ 0$ for all $\bar{\Phi}_{N,t}$ satisfying (24).
Thus all three matrices $\Sigma_{x,t}^{-,d},W_{\bar{\Phi}_{N,t}}$, and $\Omega_{\bar{\Phi}_{N,t}}$ are positive definite under (24), and therefore $r_{\mathbb{D},t}^{d}$ is a contraction for $t\geq q$. Since $r_{\mathbb{D}}$ is the corresponding time-unrolled operator, it inherits the same contraction property. ∎
C.3 Operator-norm bounds in a Bures–Wasserstein ball
We bound the operator-norm deviation between a feasible covariance matrix and its nominal counterpart when both lie in a Bures–Wasserstein ball.
Lemma 4.
Fix $\hat{X}\succ 0$. For any $X\succ 0$ with $\mathcal{B}(X,\hat{X})\leq\theta$,
$\|X-\hat{X}\|_{2}\;\leq\;\theta^{2}+2\theta\sqrt{\lambda_{\max}(\hat{X})}$.
Proof.
By Lemma˜2, there exists $U\in O(n)$ such that
$\|X^{1/2}-\hat{X}^{1/2}U\|_{2}\leq\|X^{1/2}-\hat{X}^{1/2}U\|_{F}\leq\theta.$
Let $E:=X^{1/2}-\hat{X}^{1/2}U$, so that $X^{1/2}=E+\hat{X}^{1/2}U$. Then,
$X=\hat{X}+EE^{\top}+\hat{X}^{1/2}UE^{\top}+EU^{\top}\hat{X}^{1/2},$
since $U$ is orthogonal. Taking operator norms and applying the triangle inequality and submultiplicativity of $|\cdot|_{2}$, we obtain
$\|X-\hat{X}\|_{2}\leq\|E\|_{2}^{2}+2\|\hat{X}^{1/2}\|_{2}\|E\|_{2}.$
Since $\|E\|_{2}\leq\theta$ and $\|\hat{X}^{1/2}\|_{2}=\sqrt{\lambda_{\max}(\hat{X})}$, the result follows. ∎
C.4 Proof of Theorem˜3
Proof.
From (21) and the matrix inversion lemma,
$\begin{split}\Phi_{t}&=C^{\top}\bigl(\hat{\Sigma}_{v}^{-1}-(\Sigma_{v,t}^{*})^{-1}\bigr)C+X_{t}-(I+X_{t}\Delta_{t})^{-1}X_{t}\\ &=C^{\top}(\hat{\Sigma}_{v}^{-1}-(\Sigma_{v,t}^{*})^{-1})C+X_{t}(I+\Delta_{t}X_{t})^{-1}\Delta_{t}X_{t}.\end{split}$
We bound the two terms separately.
$(i)$ Measurement uncertainty bound. Let $\delta_{v}\coloneq\|\Sigma_{v,t}^{*}-\hat{\Sigma}_{v}\|_{2}$. Since both matrices are invertible,
$\hat{\Sigma}_{v}^{-1}-(\Sigma_{v,t}^{*})^{-1}=\hat{\Sigma}_{v}^{-1}(\Sigma_{v,t}^{*}-\hat{\Sigma}_{v})(\Sigma_{v,t}^{*})^{-1}.$
By submultiplicativity, we have that
$\|\hat{\Sigma}_{v}^{-1}-(\Sigma_{v,t}^{*})^{-1}\|_{2}\leq\delta_{v}\|\hat{\Sigma}_{v}^{-1}\|_{2}\|(\Sigma_{v,t}^{*})^{-1}\|_{2}.$ (39)
On the other hand, Weyl’s inequality gives
$\lambda_{\min}(\Sigma_{v,t}^{*})\geq\lambda_{\min}(\hat{\Sigma}_{v})-\|\Sigma_{v,t}^{*}-\hat{\Sigma}_{v}\|_{2}=\lambda_{\min}(\hat{\Sigma}_{v})-\delta_{v},$
so for $\delta_{v}<\lambda_{\min}(\hat{\Sigma}_{v})$,
$\|\hat{\Sigma}_{v}^{-1}\|_{2}=\frac{1}{\lambda_{\min}(\hat{\Sigma}_{v})},$ (40)

### v3

@@ B.1 Proof of Proposition˜1
For any matrices $A,B\in\real{n\times n}$, the singular-value perturbation bounds [9, Cor. 7.3.5] state that
$\begin{split}&|\sigma_{\max}(A)-\sigma_{\max}(B)|\leq\|A-B\|_{2}\\ &|\sigma_{\min}(A)-\sigma_{\min}(B)|\leq\|A-B\|_{2}.\end{split}$ (34)
Apply (LABEL:eqn:sig_per) with $A=X^{1/2}$ and $B=\hat{X}^{1/2}U$, and combine with (33) to obtain $|\sigma_{\max}(X^{1/2})-\sigma_{\max}(\hat{X}^{1/2}U)|\leq\theta$ and $|\sigma_{\min}(\hat{X}^{1/2}U)-\sigma_{\min}(X^{1/2})|\leq\theta$. Because $U$ is orthogonal, $X\succeq 0$ and $\hat{X}\succeq 0$, we have $\sigma_{\max}(\hat{X}^{1/2}U)=\sigma_{\max}(\hat{X}^{1/2})=\sqrt{\lambda_{\max}(\hat{X})}$ and $\sigma_{\min}(\hat{X}^{1/2}U)=\sigma_{\min}(\hat{X}^{1/2})=\sqrt{\lambda_{\min}(\hat{X})}$. Substituting these into the inequalities above yields
$\begin{split}\sqrt{\lambda_{\max}(X)}&=\sigma_{\max}(X^{1/2})\leq\sqrt{\lambda_{\max}(\hat{X})}+\theta,\\ \sqrt{\lambda_{\min}(X)}&=\sigma_{\min}(X^{1/2})\geq\sqrt{\lambda_{\min}(\hat{X})}-\theta.\end{split}$
Squaring both sides and clipping the lower bound at zero yields the claim.
@@ B.2 Proof of Corollary˜2
The lower bounds follow directly from the constraints $\Sigma_{v,t}\succeq\lambda_{\min}(\hat{\Sigma}_{v,t})I_{n_{y}},\Sigma_{w,t-1}\succeq\lambda_{\min}(\hat{\Sigma}_{w,t-1})I_{n_{x}}$, and $\Sigma_{x,0}^{-}\succeq\lambda_{\min}(\hat{\Sigma}_{x,0}^{-})I_{n_{x}}$, which are imposed in (4) and (5). Thus, any maximizers $\Sigma_{v,t}^{*}$, $\Sigma_{w,t-1}^{*}$, and $\Sigma_{x,0}^{-,*}$ necessarily satisfy these inequalities, giving the claimed lower bounds.
@@ B.3 Proof of Theorem˜2
Define the posterior and prediction maps as
$\begin{split}\Psi(\Sigma_{x,t}^{-},\Sigma_{v,t})&\coloneqq\Sigma_{x,t}^{-}-\Sigma_{x,t}^{-}C_{t}^{\top}(C_{t}\Sigma_{x,t}^{-}C_{t}^{\top}+\Sigma_{v,t})^{-1}C_{t}\Sigma_{x,t}^{-}\\ \Phi(\Sigma_{x,t},\Sigma_{w,t})&\coloneqq A_{t}\Sigma_{x,t}A_{t}^{\top}+\Sigma_{w,t}.\end{split}$
The DRKF recursion is obtained by evaluating these maps at the least-favorable noise covariances: $\Sigma_{x,t}=\Psi(\Sigma_{x,t}^{-},\Sigma_{v,t}^{*})$ and $\Sigma_{x,t+1}^{-}=\Phi(\Sigma_{x,t},\Sigma_{w,t}^{*})$. For $\Sigma_{v,t}\succ 0$, the posterior covariance admits the variational representation
$\Psi(\Sigma_{x,t}^{-},\Sigma_{v,t})=\inf_{K_{t}}\Big[(I-K_{t}C_{t})\Sigma_{x,t}^{-}(I-K_{t}C_{t})^{\top}+K_{t}\Sigma_{v,t}K_{t}^{\top}\Big].$
Hence, $\Psi$ is monotone increasing in each argument: if ${\Sigma_{x,t}^{-}}^{(1)}\succeq{\Sigma_{x,t}^{-}}^{(2)}$ and $\Sigma_{v,t}^{(1)}\succeq\Sigma_{v,t}^{(2)}$, then for any $K_{t}$ the corresponding quadratic form is larger, and taking the infimum preserves the Loewner order. The map $\Phi$ is also monotone since congruence with $A_{t}$ and addition of a PSD matrix preserve the Loewner order.
Corollary˜2, together with the monotonicity of the posterior and prediction maps $\Psi$ and $\Phi$, implies by induction that, for all $t\geq 0$, the KF driven by $(\underline{\lambda}_{w,t}I_{n_{x}},\underline{\lambda}_{v,t}I_{n_{y}})$ produces covariances no larger than those of the DRKF, and the KF driven by $(\overline{\lambda}_{w,t}I_{n_{x}},\overline{\lambda}_{v,t}I_{n_{y}})$ produces covariances no smaller.
@@ Appendix C Proofs for Convergence
C.1 Proof of Lemma˜3
By Corollary˜2 and Assumption˜2, there exist constants $\underline{\lambda}_{w},\underline{\lambda}_{v}>0$ such that $\Sigma_{w,t}^{*}\succeq\underline{\lambda}_{w}I_{n_{x}}\succ 0$ and $\Sigma_{v,t}^{*}\succeq\underline{\lambda}_{v}I_{n_{y}}\succ 0$. This implies $\Sigma_{w,t}^{N,*}\succ 0$ and $\Sigma_{v,t}^{N,*}\succ 0$, and consequently,
$\Sigma_{u,t}^{N}=\mathcal{H}_{N}\Sigma_{w,t}^{N,*}\mathcal{H}_{N}^{\top}+\Sigma_{v,t}^{N,*}\succ 0.$
Using the matrix inversion lemma, the conditional covariance $\mathcal{Q}_{N,t}$ admits the equivalent representation
$\mathcal{Q}_{N,t}=\big((\Sigma_{w,t}^{N,*})^{-1}+\mathcal{H}_{N}^{\top}(\Sigma_{v,t}^{N,*})^{-1}\mathcal{H}_{N}\big)^{-1}.$
Since $(\Sigma_{w,t}^{N,*})^{-1}\succ 0$ and $\mathcal{H}_{N}^{\top}(\Sigma_{v,t}^{N,*})^{-1}\mathcal{H}_{N}\succeq 0$, it follows that $\mathcal{Q}_{N,t}\succ 0$. Moreover, $\mathcal{R}_{N}$ has full row rank for all $N\geq 1$. Hence, for any nonzero $x\in\real{n_{x}}$,
$x^{\top}W_{N,t}x=(\mathcal{R}_{N}^{\top}x)^{\top}\mathcal{Q}_{N,t}(\mathcal{R}_{N}^{\top}x)>0,$
which proves that $W_{N,t}\succ 0$.
Finally, since $\Sigma_{u,t}^{N}\succ 0$, we have $(\Sigma_{u,t}^{N})^{-1}\succ 0$. By Assumption˜3, the observability of $(A,C)$ implies that $\mathcal{O}_{N}$ has full column rank for any $N\geq n_{x}$. Therefore, for any nonzero $x\in\real{n_{x}}$,
$x^{\top}\Omega_{N,t}x=(\mathcal{O}_{N}x)^{\top}(\Sigma_{u,t}^{N})^{-1}(\mathcal{O}_{N}x)>0,$
which establishes $\Omega_{N,t}\succ 0$ for all $N\geq n_{x}$.
@@ C.2 Proof of Proposition˜2
Fix $\theta_{w}$ and $\theta_{v}$ and let $\underline{\lambda}_{w}\coloneq\lambda_{\min}(\hat{\Sigma}_{w})$, $\underline{\lambda}_{v}\coloneq\lambda_{\min}(\hat{\Sigma}_{v})$, $\overline{\lambda}_{w}\coloneq\overline{\lambda}(\hat{\Sigma}_{w},\theta_{w})$, and $\overline{\lambda}_{v}\coloneq\overline{\lambda}(\hat{\Sigma}_{v},\theta_{v})$ as in Corollary˜2. Then, the least-favorable covariances satisfy $\underline{\lambda}_{w}I\preceq\Sigma_{w,t}^{N,*}\preceq\overline{\lambda}_{w}I$ and $\underline{\lambda}_{v}I\preceq\Sigma_{v,t}^{N,*}\preceq\overline{\lambda}_{v}I$. Hence,
$\Sigma_{u,t}^{N}=\mathcal{H}_{N}\Sigma_{w,t}^{N,*}\mathcal{H}_{N}^{\top}+\Sigma_{v,t}^{N,*}\preceq(\overline{\lambda}_{w}\|\mathcal{H}_{N}\|_{2}^{2}+\overline{\lambda}_{v})I,$
and $\Sigma_{u,t}^{N}\succeq\underline{\lambda}_{v}I$. Therefore,
$\lambda_{\min}(\Omega_{N,t})\geq\frac{\sigma_{\min}(\mathcal{O}_{N})^{2}}{\lambda_{\max}(\Sigma_{u,t}^{N})}\geq\frac{\sigma_{\min}(\mathcal{O}_{N})^{2}}{\overline{\lambda}_{v}+\overline{\lambda}_{w}\|\mathcal{H}_{N}\|_{2}^{2}},$
which implies
$\|\Omega_{N,t}^{-1}\|_{2}\leq\frac{\overline{\lambda}_{v}+\overline{\lambda}_{w}\|\mathcal{H}_{N}\|_{2}^{2}}{\sigma_{\min}(\mathcal{O}_{N})^{2}}.$
Next, using the definition of $\mathcal{G}_{N,t}$, we obtain
$\|\mathcal{G}_{N,t}\|_{2}\leq\|\Sigma_{w,t}^{N,*}\|_{2}\|\mathcal{H}_{N}\|_{2}\,\|(\Sigma_{u,t}^{N})^{-1}\|_{2}\leq\frac{\overline{\lambda}_{w}}{\underline{\lambda}_{v}}\|\mathcal{H}_{N}\|_{2}.$
Consequently,
$\begin{split}\|\alpha_{N,t}\|_{2}&\leq\|A^{N}\|_{2}+\|\mathcal{R}_{N}\|_{2}\|\mathcal{G}_{N,t}\|_{2}\|\mathcal{O}_{N}\|_{2}\\ &\leq\|A^{N}\|_{2}+\frac{\overline{\lambda}_{w}}{\underline{\lambda}_{v}}\|\mathcal{R}_{N}\|_{2}\|\mathcal{H}_{N}\|_{2}\|\mathcal{O}_{N}\|_{2}.\end{split}$
Finally, we have that
$\begin{split}\lambda_{\min}(\mathcal{Q}_{N,t})&\geq\frac{1}{\lambda_{\max}((\Sigma_{w,t}^{N,*})^{-1})+\lambda_{\max}(\mathcal{H}_{N}^{\top}(\Sigma_{v,t}^{N,*})^{-1}\mathcal{H}_{N})}\\ &\geq\frac{1}{1/\underline{\lambda}_{w}+\|\mathcal{H}_{N}\|_{2}^{2}/\underline{\lambda}_{v}}.\end{split}$
Since $\mathcal{R}_{N}\mathcal{R}_{N}^{\top}\succeq I$, we have $\lambda_{\min}(W_{N,t})\geq\lambda_{\min}(\mathcal{Q}_{N,t})$, and therefore
$\|W_{N,t}^{-1}\|_{2}\leq 1/\underline{\lambda}_{w}+\|\mathcal{H}_{N}\|_{2}^{2}/\underline{\lambda}_{v}.$
Combining the above bounds yields
$\|M_{N,t}\|_{2}\leq\|\Omega_{N,t}^{-1}\|_{2}\|\alpha_{N,t}\|_{2}^{2}\|W_{N,t}^{-1}\|_{2}\leq\kappa_{N}(\theta_{v},\theta_{w}).$
Substituting into (24) and using the monotonicity of $s\mapsto(\frac{\sqrt{s}}{1+\sqrt{1+s}})^{2}$ on $s\geq 0$ gives $\xi_{N,t}\leq\bar{\xi}_{N}(\theta_{v},\theta_{w})$; inequality (26) then ensures $\xi_{N,t}\leq\rho$ for all $t\geq 0$.
C.3 Proof of Theorem˜3
Fix $t\geq 0$ and $N\geq n_{x}$. By Lemma˜3, we have $W_{N,t}\succ 0$ and $\Omega_{N,t}\succ 0$. Hence, the downsampled DR Riccati mapping $r_{\mathbb{D},t}^{d}$ in (22) is well-defined on $\mathbb{S}^{n_{x}}_{++}$ and has the robust Riccati form considered in [14, Thm. 5.3]. Therefore, $r_{\mathbb{D},t}^{d}$ is a strict contraction with respect to Thompson’s part metric on $\mathbb{S}^{n_{x}}_{++}$, with contraction factor $\xi_{N,t}\in(0,1)$ satisfying the bound (24).
By Proposition˜2, we have $\xi_{N,t}\leq\bar{\xi}_{N}(\theta_{v},\theta_{w})<1$ for all $t\geq 0$, and thus $r_{\mathbb{D},t}^{d}$ is uniformly contractive. Under Assumption˜2, the one-step DR Riccati update is stationary, so the downsampled map coincides with the NN-fold iterate of a single time-invariant map. The contraction argument for downsampled iterations in [36] then implies that the one-step DR Riccati recursion (19) converges to a unique fixed point $\Sigma_{x,\infty}^{-}\succ 0$.
Finally, although the stage-wise SDP may admit multiple optimizers in $\Sigma_{v,t}^{*}$, the corresponding DR Kalman gain is unique. By Lemma˜1, for a fixed $\Sigma_{x,t}^{-}$ the stage-wise DRSE problem amounts to minimizing over $K$ the worst-case posterior MSE, with the adversary maximizing over $\Sigma_{v,t}$ in the SDP’s feasible set, denoted by $\mathcal{A}_{v,t}$. Using the standard KF covariance identity, the resulting worst-case one-step MSE is
$\sup_{\Sigma_{v,t}\in\mathcal{A}_{v,t}}\operatorname{Tr}\big[(I-KC)\Sigma_{x,t}^{-}(I-KC)^{\top}+K\Sigma_{v,t}K^{\top}\big].$
For each feasible $\Sigma_{v,t}$, the associated quadratic map
$K\mapsto\operatorname{Tr}\big[(I-KC)\Sigma_{x,t}^{-}(I-KC)^{\top}+K\Sigma_{v,t}K^{\top}\big]$
is $\underline{\lambda}_{v}$-strongly convex in $K$ (with respect to $\|\cdot\|_{F}$). Taking the supremum over the compact feasible set $\mathcal{A}_{v,t}$ preserves strong convexity, and hence the minimizer $K_{t}$ is unique. Moreover, this worst-case objective is continuous in $(\Sigma_{x,t}^{-},K)$, so the unique minimizer $K_{t}$ depends continuously on $\Sigma_{x,t}^{-}$. Therefore, $\Sigma_{x,t}^{-}\to\Sigma_{x,\infty}^{-}$ implies that the sequence $\{K_{t}\}$ converges.
C.4 Proof of Corollary˜3
For each $t\geq 1$, let $z_{t}:=(\Sigma_{x,t}^{-},\Sigma_{x,t},\Sigma_{w,t-1}^{*},\Sigma_{v,t}^{*},Y_{t},Z_{t})$ denote a primal optimal solution of the stage-wise SDP (7). By Theorem˜3, $(\Sigma_{x,t}^{-},\Sigma_{x,t})\to(\Sigma_{x,\infty}^{-},\Sigma_{x,\infty})$.
Using the LMIs $\begin{bmatrix}\hat{\Sigma}_{w}&Y_{t}\\ Y_{t}^{\top}&\Sigma_{w,t-1}^{*}\end{bmatrix}\succeq 0$ and $\begin{bmatrix}\hat{\Sigma}_{v}&Z_{t}\\ Z_{t}^{\top}&\Sigma_{v,t}^{*}\end{bmatrix}\succeq 0$, together with $\hat{\Sigma}_{w}\succ 0$, $\hat{\Sigma}_{v}\succ 0$, Schur complements yield $\Sigma_{w,t-1}^{*}\succeq Y_{t}^{\top}\hat{\Sigma}_{w}^{-1}Y_{t}$ and $\Sigma_{v,t}^{*}\succeq Z_{t}^{\top}\hat{\Sigma}_{v}^{-1}Z_{t}$. Hence, it follows from Corollary˜2 that
$\|Y_{t}\|_{2}^{2}\leq\|\hat{\Sigma}_{w}\|_{2}\ \|\Sigma_{w,t-1}^{*}\|_{2}\leq\|\hat{\Sigma}_{w}\|_{2}\ \bar{\lambda}_{w}$